OpenStudy (anonymous):

Another calculus problem.. >_< Use the substitution x=8tan() to evaluate the indefinite integral: 72/(x^2 sqrt(x^2+64)) Thanks! :D

OpenStudy (anonymous):

waht is 8tan(), do you ean for theta?

OpenStudy (anonymous):

*mean

OpenStudy (anonymous):

Yeah...

OpenStudy (anonymous):

hint* if x = 8tan() then x^2 = 64tan^2() and remember tan^2() + 1 = Sec^2()

OpenStudy (anonymous):

so first try plugging 64tan^2() for all x^2

OpenStudy (anonymous):

Yeah.. That's what I did.. But I got stuck after that.. When I have \[(9/64)\int\limits 1/(\tan^2\theta \sec \theta)\]

OpenStudy (anonymous):

remember your indefinite integral needs a (dx) if x = 8tan() then dx = 8sec^2()d()

OpenStudy (anonymous):

\[\int\limits_{}^{}(72/x^2\sqrt{x^2 + 64} )dx = 9/8\int\limits_{}^{}[/\tan^2\theta \sec \theta]\sec^2\theta\]

OpenStudy (anonymous):

\[ 9/8\int\limits_{}^{}[1/(\tan^2\theta)(\sec \theta)] \sec^2\theta d \theta = 9/8\int\limits_{}^{}(\sec \theta / \tan^2\theta) d \theta\] \[= 9/8\int\limits_{}^{}\csc \theta \cot \theta d \theta\] \[= -9/8 \csc \theta\]

OpenStudy (anonymous):

Ohh.. That makes sense! Thank you so much! :D

OpenStudy (anonymous):

haha i remember one of my friends said " you don't need trig for calculus". He didnt pass calc 2

OpenStudy (anonymous):

Ohh wow.. That's really funny.. >_<